# Statistical Guarantees in Synthetic Data through Conformal Adversarial Generation

### v1

@@ I Introduction
Existing approaches for synthetic data generation include GANs [1], Variational Autoencoders (VAEs) [2], Diffusion Models [3], and Normalizing Flows [4]. While these methods have demonstrated impressive capabilities in generating high-quality samples, they suffer from significant limitations. GANs are prone to mode collapse and training instability [5]. VAEs often produce blurry outputs and struggle with complex distributions [6]. Diffusion models, despite recent advances [7], require lengthy sampling processes. Most critically, none of these approaches provide rigorous guarantees regarding the statistical properties of the generated samples [8], making their outputs unreliable for high-stakes applications in healthcare [9], finance [10], and autonomous systems [11]. The motivation for conformalized GANs (C-GANs) stems from the need for statistically valid synthetic data with quantifiable uncertainty. By conformally calibrating both the generator and discriminator components, we enhance the quality and reliability of synthetic datasets. This approach provides theoretical guarantees about the statistical fidelity of generated samples relative to the real data distribution, addressing a critical gap in existing generative models [12].
@@ II Motivation
The imperative to develop conformalized GANs arises from several interconnected challenges in contemporary generative modeling. First, the absence of rigorous statistical guarantees in existing generative models represents a fundamental impediment to their deployment in high-stakes applications [13]. In clinical settings, for instance, synthetic patient data lacking verifiable statistical properties may lead to erroneous conclusions with potentially severe consequences [14]. Financial institutions similarly require synthetic market scenarios with quantifiable uncertainty for robust risk assessment and regulatory compliance [15].
@@ VIII-D Limitations and Future Directions
Fourth, alternative metrics that we did not explore in this investigation include privacy preservation guarantees and adversarial robustness. Future work should examine whether conformalized GANs offer inherent advantages in privacy-preserving synthetic data generation [23] and resistance to adversarial attacks [24], both increasingly critical considerations in real-world deployments.

### v3

@@ I Introduction
Existing approaches for synthetic data generation include GANs [1], Variational Autoencoders (VAEs) [2], Diffusion Models [3], and Normalizing Flows [4]. While these methods have demonstrated impressive capabilities in generating high-quality samples, they suffer from significant limitations. GANs are prone to mode collapse and training instability [5]. VAEs often produce blurry outputs and struggle with complex distributions [6]. Diffusion models, despite recent advances [7], require lengthy sampling processes. Most critically, none of these approaches provide rigorous guarantees regarding the statistical properties of the generated samples [8], making their outputs unreliable for high-stakes applications in healthcare [9], finance [10], and autonomous systems [11]. The motivation for conformalized GANs (C-GANs) stems from the need for statistically valid synthetic data with quantifiable uncertainty. By conformally calibrating both the generator and discriminator components, we enhance the quality and reliability of synthetic datasets. This approach provides theoretical guarantees about the statistical fidelity of generated samples relative to the real data distribution, addressing a critical gap in existing generative models [12] [13].
@@ II Motivation
The imperative to develop conformalized GANs arises from several interconnected challenges in contemporary generative modeling. First, the absence of rigorous statistical guarantees in existing generative models represents a fundamental impediment to their deployment in high-stakes applications [14] [15]. In clinical settings, for instance, synthetic patient data lacking verifiable statistical properties may lead to erroneous conclusions with potentially severe consequences [16] [17]. Financial institutions similarly require synthetic market scenarios with quantifiable uncertainty for robust risk assessment and regulatory compliance [18].
@@ VIII-D Limitations and Future Directions
Fourth, alternative metrics that we did not explore in this investigation include privacy preservation guarantees and adversarial robustness. Future work should examine whether conformalized GANs offer inherent advantages in privacy-preserving synthetic data generation [26, 27] and resistance to adversarial attacks [28], both increasingly critical considerations in real-world deployments.